8. The average drive to work is 9.6 miles. Assume standard deviation is 1.8. If random sample of 36 employed who derive to work are selected, find the probability that mean of sample miles driven to work is between 9 and 10 miles.


1
Expert's answer
2022-06-15T13:06:55-0400

Solution:

Let's denote given values:

"\mu=9.6 miles" - population mean;

"\sigma=1.8 miles" -standard deviation;

"n=36" - sample number;

"X_1=9miles;" "X_2=10miles" -mean samples.

Find "z:"

"z_1=\frac{X_1-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{9-9.6}{\frac{1.8}{\sqrt{36}}}=-2;"

"z_1= -2;" so, probability from z table "0.028" , it means "2.8\%" of X less than 9;

Find "z_2:"

"z_2=\frac{X_2-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{10-9.6}{\frac{1.8}{\sqrt{36}}}=1.33;"

From z table: "z_2=1.33," so probability "0.0918". "1-0.918=0.9082;" "90.82\%" of X less than 10. So,

"P(9<X<10)=0.9082-0.028=0.8802;"

Answer:

"P(9<X<10)=88.02\%."